Phyllis bought 14 yards of material to make chair cushions. She cut the material into pieces 1 3/4 yards long to make each cushion. How many cushions did Phyllis make?
Answer by stanbon(75887)

Phyllis bought 14 yards of material to make chair cushions. She cut the material into pieces 1 3/4 yards long to make each cushion. How many cushions did Phyllis make?
-------------------
# of cushions = 14/(1 3/4) = 14/(7/4) = 14*(4/7) = 2*4 = 8
================
Cheers,
Stan H.
==============